The fundamental frequency of a closed pipe is 293 Hz when the air in it is at a temperature of 20°C. What will be its fundamental frequency when the temperature changes to 22°C?

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. a) Here given, L base r = 1.0/2 = 0.5 m, d base a = 6.5 cm = 6.5 × 10^–2 m As Kundt’s tube apparatus is a closed organ pipe, its fundamental frequency ⇒n = V base r/4l base r ⇒ V base r = 2600 * 4 * 0.5 = 5200 m/s. b) V base r/V base a = 2L base r/d base a ⇒ v base a = 5200 * 6.5 * 10^-2/2 * 0.5 = 338 m/s.